# On metric choice in dimension reduction for Fréchet regression

### v1

@@ 6 Discussion
A future study that explores how metric choice affects the sufficient dimension reductions estimates in regression settings where both the predictor and response to belong to random metric spaces will be a reasonable extension.
[…]
@@
Proof .3 (of Proposition 2.4).
$Y_{1}\succeq_{1}Y_{2}$ if and only if $F_{Y_{1}}(t)\leq F_{Y_{2}}(t),\forall t\in[0,\infty)$. Thus, for any non-decreasing function $g(y)$, we have $\mathrm{E}_{Y_{1}}(g(y))\geq\mathrm{E}_{Y_{2}}(g(y))$. Therefore, $W_{1}(Y_{1},Y_{2})=\mathrm{E}|Y_{1}-Y_{2}|=\mathrm{E}[\mathrm{E}|Y_{1}-Y_{2}|\mid Y_{2}<Y_{1},Y_{2}]=\mathrm{E}(\mathrm{E}(Y_{1}-Y_{2}\mid Y_{2}<Y_{1},Y_{2}))=\mathrm{E}(Y_{1})-\mathrm{E}(Y_{2})$.
Proof .4 (of Proposition 2.5).
@@
Now, suppose $\widetilde{Y}_{\rho_{2}}=h_{1}(\bm{\eta}_{1}^{\top}{\bf X})+h_{2}(\bm{\eta}_{2}^{\top}{\bf X})\epsilon$, where $\bm{\eta}=(\bm{\eta}_{1},\bm{\eta}_{2})\in\mathbb{R}^{d}$ is such that $\bm{\eta}^{\top}\bm{\eta}={\bf I}_{d}$. Then $\mathcal{S}_{\mathrm{E}(\widetilde{Y}_{\rho_{2}}|{\bf X})}=\mathrm{span}(\bm{\eta}_{1})$ while $\mathcal{S}_{\widetilde{Y}_{\rho_{2}}|{\bf X}}=\mathrm{span}(\bm{\eta}_{1},\bm{\eta}_{2})$. Therefore, estimating $\mathcal{S}_{\mathrm{E}(\widetilde{Y}_{\rho_{2}}|{\bf X})}$ leads to some loss of regression information.
Some common distributional distances
Hellinger Distance between two distributions

### v2

@@ 6 Discussion
The conflicting conclusions from the ${\mathcal{S}}_{\mathrm{E}(Y|{\bf X})}$ and ${\mathcal{S}}_{Y|{\bf X}}$ estimates based on different metrics buttresses the point that metric choice plays a crucial role, especially when dealing with the mean space, as we observed in the numerical studies. A future study that explores how metric choice affects the sufficient dimension reductions estimates in regression settings where both the predictor and response to belong to random metric spaces will be a reasonable extension.
@@
Proof of Proposition 3.
$Y_{1}\succeq_{1}Y_{2}$ if and only if $F_{Y_{1}}(t)\leq F_{Y_{2}}(t),\forall t\in[0,\infty)$. Thus, for any non-decreasing function $g(y)$, we have $\mathrm{E}_{Y_{1}}(g(y))\geq\mathrm{E}_{Y_{2}}(g(y))$. Therefore, $W_{1}(Y_{1},Y_{2})=\mathrm{E}|Y_{1}-Y_{2}|=\mathrm{E}[\mathrm{E}|Y_{1}-Y_{2}|\mid Y_{2}<Y_{1},Y_{2}]=\mathrm{E}(\mathrm{E}(Y_{1}-Y_{2}\mid Y_{2}<Y_{1},Y_{2}))=\mathrm{E}(Y_{1})-\mathrm{E}(Y_{2})$. ∎
Proof of Proposition 4.
@@
Now, suppose $\widetilde{Y}_{\rho_{2}}=h_{1}(\bm{\eta}_{1}^{\top}{\bf X})+h_{2}(\bm{\eta}_{2}^{\top}{\bf X})\epsilon$, where $\bm{\eta}=(\bm{\eta}_{1},\bm{\eta}_{2})\in\mathbb{R}^{d}$ is such that $\bm{\eta}^{\top}\bm{\eta}={\bf I}_{d}$. Then $\mathcal{S}_{\mathrm{E}(\widetilde{Y}_{\rho_{2}}|{\bf X})}=\mathrm{span}(\bm{\eta}_{1})$ while $\mathcal{S}_{\widetilde{Y}_{\rho_{2}}|{\bf X}}=\mathrm{span}(\bm{\eta}_{1},\bm{\eta}_{2})$. Therefore, estimating $\mathcal{S}_{\mathrm{E}(\widetilde{Y}_{\rho_{2}}|{\bf X})}$ leads to some loss of regression information. ∎
Hellinger distances